# The Thurston norm of graph manifolds

### v1

@@ 1 Introduction
The Thurston norm [Th86] is a powerful tool for analyzing the topology of a 3-manifold by studying embedded surfaces sitting inside the $3$-manifold. Given a compact orientable 3-manifold $M$, the Thurston norm is a seminorm $x$ defined on the vector space $H_{2}(M,\partial M;\operatorname{\mathbb{R}})$. If $S$ is some properly embedded oriented surface in $M$, then $-x([S])$ is the maximal Euler characteristic among the properly embedded oriented surfaces homologous to $S$, computed after discarding any component being a sphere, a disc, a torus or an annulus. For instance, in a knot complement, the Thurston norm detects the minimal genus of the knot.
The unit ball $B_{x}$ of $x$ is a finite (possibly unbounded) convex polyhedron (cf. [Th86]), thus the whole information carried by $x$ is determined by a finite amount of data, namely the position of the vertices of $B_{x}$.
Unfortunately, the Thurston norm has been described just in some special families of manifolds (see for example [Th86, Li08, OS08, PSV19, Ch23, Ci24]). In [Ci24], we focused on the manifolds obtained as $2$-bridge link complements. There, the Thurston norm is a seminorm on the plane and happens to be always quite simple: the unit polygon has at most $8$ vertices, lying along the axes and the principal bisectors. However, by means of this result, we showed that the Thurston norms obtainable on the plane can have unit balls with an arbitrary number of sides, even if we restrict to complements in $S^{3}$ of links with two unknotted components.
In the present paper, we move our attention to a different family of $3$-manifolds, namely graph manifolds. In many cases, the second homology of a graph manifold has dimension $\geq 3$, a setting where even less is known about the possible shapes of Thurston unit polyhedra. The advantage of graph manifolds is that their surfaces can be easily understood. Indeed, a graph manifold decomposes into Seifert fibered pieces, and an incompressible and boundary-incompressible surface in a Seifert fibered manifold can be isotoped to become either horizontal or vertical. Moreover, a useful result by Neumann [Ne97] helps us to understand when surfaces inside the different pieces can patch together to give a properly embedded surface in the ambient graph manifold.
@@
Theorem 4.1. Let $\beta_{1},...,\beta_{n}\in\operatorname{\mathbb{Q}}^{d}$ be such that the function
@@
is a norm on $\operatorname{\mathbb{R}}^{d}$, and let $g_{1},...,g_{n}$ be nonnegative integers. There exist a closed oriented good graph manifold $M$ with base surfaces of genera $g_{1},...,g_{n}$ and a rational basis on $H_{2}(M)$, such that the nonvanishing Thurston norm on $H_{2}^{nv}(M;\operatorname{\mathbb{R}})$ in the induced basis coincides with |||||$\cdot$||| | ⋅ | |. Moreover, we can ask for either all or none of the top-dimensional Thurston cones of $H_{2}(M)$ to be fibered.
In Theorem 4.1, good graph manifold means that its Seifert fibered JSJ pieces all have orientable base orbifold, whereas the nonvanishing homology $H_{2}^{nv}(M\operatorname{\mathbb{R}})$ and the nonvanishing Thurston norm refer to the quotient of $H_{2}(M;\operatorname{\mathbb{R}})$ by the null-space of the Thurston norm and the induced norm respectively. Notice that, in the case of good graph manifolds, we can also compute the dimension of the null-space of the Thurston norm.
@@
Observe that, in [Th86], Thurston has shown a result that is stronger than Corollary 4.4 in the way that it just uses integral bases and realises the polygons directly on $H_{2}(M;\operatorname{\mathbb{R}})$. At the same time, the new feature of Corollary 4.4 is that we can keep track of more structure on the manifolds used to realise the given planar norms, namely their being good graph manifolds and whether they fiber over the circle or not.
@@
Corollary 5.8 Let $P\subset\operatorname{\mathbb{R}}^{n}$ be a polyhedron with rational vertices which is symmetric through the origin. There is a closed oriented good graph manifold $M$, such that there is a rational basis in which $H_{2}^{nv}(M)\cong\operatorname{\mathbb{R}}^{n}$ and the partition of $H_{2}^{nv}(M)$ into cones over the faces of the nonvanishing Thurston unit ball is a refinement of the partition of $\operatorname{\mathbb{R}}^{n}$ into cones over the faces of $\partial P$. Moreover, we can ask for either all or none of the top-dimensional Thurston cones of $H_{2}(M)$ to be fibered.
Structure of the paper. In Section 2 we recall the definitions and basic properties of the main objects on focus: the Thurston norm, Seifert fibered manifolds, and graph manifolds. Along the way, we introduce a result by Neumann [Ne97] that facilitates our understanding of the surfaces sitting inside graph manifolds. In Section 3 we compute the norm of a graph manifold. In Section 4, a characterization of all norms that can be obtained as nonvanishing Thurston norm of graph manifolds is achieved (see Theorem 4.1). Finally, in Section 5 we focus on the polyhedra arising as unit balls of Thurston norms on graph manifolds. This last section aims to understand the geometry of such polyhedra and their properties better. In this way, we will have an idea of how “generic” such polyhedra are, inside the class of all polyhedra or the class of those polyhedra arising as Thurston unit balls of some $3$-manifold.
Aknowledgements. I would like to thank my PhD advisor Mehdi Yazdi for useful discussions on the topic and for his supervision.
@@ 2 Preliminaries
We now recall the basic definitions and properties of the Thurston norm. For further details, we refer to [Th86]. From now on, $M$ will always be a $3$-dimensional compact orientable smooth manifold.
@@
[Th86] If $S\subset M$ is the fiber of a fibration of $M$ over $S^{1}$, then $[S]\in H_{2}(M,\partial M)$ lies in the interior of the cone over a top-dimensional boundary face of the Thurston ball $B_{x}$. Viceversa, there is a collection of top-dimensional boundary faces of $B_{x}$ such that every integral class lying in the interior of the cone over one such face is represented by a fiber of a fibration of $M$ over $S^{1}$.
@@
[Lack21, Proposition 11.2] Let $M$ be a compact oriented $3$-manifold with incompressible boundary. Let $T\subset M$ be a disjoint union of incompressible tori and let $M^{\prime}$ be $M$ cut along $T$, seen as a submanifold of $M$. Consider a properly embedded oriented surface $S\subset M$, being transverse to $T$. Indicate with $x_{M}$ and $x_{M^{\prime}}$ the Thurston norms on $H_{2}(M,\partial M)$ and $H_{2}(M^{\prime},\partial M^{\prime})$ respectively. The following equality holds
@@ 2.2 Seifert fibered manifolds and their Thurston norm
We now introduce the building blocks for graph manifolds, which are Seifert fibered manifolds. For further details, see [Mar] for example.
@@
The manifold $(\Sigma,(p_{1},q_{1}),...,(p_{r},q_{r}))$ can be endowed with a singular fibration by circles (see [Mar]) over the base orbifold $(\Sigma,p_{1},...,p_{r})$, i.e. the orbifold whose base surface is $\Sigma$ with $r$ cone points, with orders $p_{1},...,p_{r}$. This singular fibration is restricted to the regular circle bundle on $\Sigma\widetilde{\times}S^{1}-\left(\cup_{i}N(c_{i})\right)$ and has singular fibers only corresponding to the cone points of $(\Sigma,p_{1},...,p_{r})$. This special type of singular fibration is called Seifert fibration, hence we call $(\Sigma,(p_{1},q_{1}),...,(p_{r},q_{r}))$ a Seifert fibered manifold (over the base orbifold $(\Sigma,p_{1},...,p_{r})$).
The classification of Seifert fibered manifolds up to homeomorphism is completely understood [Mar]. A weaker problem concerns the classification of Seifert fibered manifolds up to isomorphism. If we are given two Seifert fibered manifolds $M$ and $M^{\prime}$ with fixed Seifert fibrations on them, we say they are isomorphic if there is a diffeomorphism between $M$ and $M^{\prime}$ respecting the fibrations. The same manifold can admit two different (non-isomorphic) Seifert fibrations.
@@
[Mar, Theorem 10.4.19] If $\partial\Sigma\neq\emptyset$ and $\Sigma$ is not a disc nor a Moebius band, then $(\Sigma,(p_{1},q_{1}),...,(p_{r},q_{r}))$ admits only one Seifert fibration up to isomorphism.
@@
When dealing with Seifert fibered manifolds, two important invariants are always concerned, namely the base orbifold Euler characteristic and the Euler number. Here, we will use the conventions of [Ne97].
@@
For the remaining part of this section, let $M=(\Sigma,(p_{1},q_{1}),...,(p_{r},q_{r}))$ with the induced Seifert fibration.
The fibration in $M$ helps us to describe the surfaces sitting inside $M$.
@@
Let $S\subset M$ be a properly embedded surface. We say that $S$ is vertical if it is a union of nonsingular circle fibers. We say that $S$ is horizontal if it intersects the fibers transversely.
@@
[Mar, Propositions 10.4.9 and 10.4.10] Let $S\subset M$ be a properly embedded surface and suppose $M$ is not diffeomorphic to $S^{2}\times S^{1}$ nor to $\mathbb{RP}^{2}\widetilde{\times}S^{1}$. If $S$ is incompressible, $\partial$-incompressible and no component of $S$ is $\partial$-parallel, then $S$ is isotopic to either a vertical or a horizontal surface.
@@
Let $S\subset M$ be a properly embedded connected and oriented surface. Up to homology, we can assume $S$ is incompressible and $\partial$-incompressible. In particular, after possibly discarding null-homologous components, we can suppose $S$ is either vertical or horizontal.
@@
As proved in [Mar, Proposition 10.4.8], if $M$ is closed with $e(M)\neq 0$, then $M$ contains no horizontal surface, so $H_{2}^{nv}(M,\partial M)=0$.
Suppose that $S$ is horizontal. Then, the projection on $\Sigma$ is a degree-$d$ orbifold covering, where $d$ is the number of times that $S$ transversely intersects each fiber. Thus,
@@
By cutting $M$ along $S$, we obtain an oriented interval bundle over $S$, hence $S\times[0,1]$. So, the manifold $M$ fibers over the circle with fiber $S$, and this implies that $S$ is norm-minimising in $H_{2}(M,\partial M)$ [Th86], i.e.
$x([S])=\chi_{-}(S)=d\max(0,-\chi^{orb}(M)).$
Finally, suppose $S^{\prime}$ is another oriented horizontal surface, transversely intersecting each fiber $d^{\prime}$ times. After possibly reversing the orientation of $S^{\prime}$, we can suppose that the surface given by the oriented cut-and-paste $\overline{S}$ of $d^{\prime}$ parallel copies of $S$ and $d$ parallel copies of $S^{\prime}$ has algebraic intersection 00 with each regular fiber. It follows that $\overline{S}$ is homologous to a vertical surface, thus annihilating the Thurston norm. This proves that the nonvanishing homology has dimension at most one, and the reasoning above shows exactly when it vanishes. ∎
@@ 2.3 Graph manifolds and their reduced plumbing matrices
In this section, we introduce graph manifolds and some results by Neumann [Ne97] that will be essential for our ends. We will be interested only in the closed orientable case, even if the stated facts have broader generality.
@@
When $M$ is a graph manifold, there is a canonical family of decomposing tori in $M$ up to isotopy. This family is characterized as the minimal one (with respect to the inclusion) cutting $M$ into Seifert fibered pieces, and the decomposition is the JSJ decomposition of $M$ [Mar]. When working with a graph manifold $M$ and unless otherwise mentioned, we will always consider its subdivision into Seifert fibered pieces coming from the JSJ decomposition.
@@
[Ne97, Lemma 4.2] Let $M$ be closed oriented good graph manifold with pieces $M_{1},...,M_{n}$. Assume that $M$ is not a torus-bundle over $S^{1}$. Let $(l_{1},...,l_{n})$ be a tuple of integers. The following properties are equivalent:
  • There are a natural $N>0$ and a properly embedded oriented surface $S\subset M$ intersecting the regular fibers of $M_{i}$ algebraically $Nl_{i}$ times, for each $i=1,...,n$;
@@
In the original version of [Ne97, Lemma 4.2], in the place of the first property there was
@@
Observe that the corrected version of Lemma 2.17 is enough to prove the results of [Ne97], by just replacing the original lemma.
@@
Sketch of the proof. In [Ne97], Neumann considers $M$ as a plumbing manifold, often relying on the results of [Ne81]. A plumbing manifold is a $3$-manifold $N$ that admits a family of decomposing tori $\mathcal{T}$ such that:
@@
The base surfaces of the plumbing pieces $M_{1}^{\prime},...,M_{k}^{\prime}$ of $M$ give a basis of $H_{2}(X;\operatorname{\mathbb{Z}})$. In such basis, $\alpha$ sends the class of a properly embedded oriented surface $S\subset M$ to the tuple $(\ell_{1}^{\prime},...,\ell_{k}^{\prime})$, where $\ell_{i}^{\prime}$ is the algebraic intersection of $S$ with the fiber of $M_{i}^{\prime}$. The map $\beta$ is the intersection form, represented by a matrix $\mathcal{S}$ in the aforementioned basis. Now, we just need to establish the relation between the tuples $(\ell_{1}^{\prime},...,\ell_{k}^{\prime})$ and $(\ell_{1},...,\ell_{n})$, as well as the relation between $\mathcal{S}$ and the reducing plumbing matrix $A$ of $M$.
Each piece $M_{j}$ of the JSJ decomposition of $M$ corresponds to a subgraph $\Gamma_{j}$ of $\Gamma^{\prime}$, and the subgraphs obtained in this way all together cover $\Gamma^{\prime}$, while pairwise intersecting just in a union of vertices. In every subgraph $\Gamma_{j}$, there is a preferred vertex, called the node of the subgraph. The node is a vertex with positive genus or with degree $\geq 3$. Label the vertices of $\Gamma^{\prime}$ so that the nodes are labelled $1,...,n$. The matrix $\mathcal{S}$ can be modified through a sequence of column and row moves to get the direct sum of $A$ with a diagonal $(k-n)\times(k-n)$ matrix. Each row (resp. column) move consists of adding to a row (resp. column) a rational multiple of one of the last $k-n$ rows (resp. columns). In particular, if $(\ell_{1}^{\prime},...,\ell_{k}^{\prime})\in\operatorname{\mathbb{Z}}^{k}$ annihilates $\mathcal{S}$, then $(\ell_{1}^{\prime},...,\ell_{n}^{\prime})$ annihilates $A$. Viceversa, if $(\ell_{1},...,\ell_{n})\in\operatorname{\mathbb{Z}}^{n}$ annihilates $A$, then there are $\ell_{n+1}^{\prime},...,\ell_{k}^{\prime}\in\operatorname{\mathbb{Q}}$ such that $(\ell_{1},...,\ell_{n},\ell_{n+1}^{\prime},...,\ell_{k}^{\prime})$ annihilates $\mathcal{S}$. Since $\ell_{n+1}^{\prime},...,\ell_{k}^{\prime}$ can be nonintegral, we could need to multiply $(\ell_{1},...,\ell_{n},\ell_{n+1}^{\prime},...,\ell_{k}^{\prime})$ by a nonzero natural number to obtain a tuple of integers annihilating $\mathcal{S}$.

### v2

@@ 1 Introduction
The Thurston norm [13] is a powerful tool for analyzing the topology of a 3-manifold by studying embedded surfaces sitting inside the $3$-manifold. Given a compact orientable 3-manifold $M$, the Thurston norm is a seminorm $x$ defined on the vector space $H_{2}(M,\partial M;\operatorname{\mathbb{R}})$. If $S$ is some properly embedded oriented surface in $M$, then $-x([S])$ is the maximal Euler characteristic among the properly embedded oriented surfaces homologous to $S$, computed after discarding any component being a sphere, a disc, a torus or an annulus. For instance, in a knot complement, the Thurston norm detects the minimal genus of the knot.
The unit ball $B_{x}$ of $x$ is a finite (possibly unbounded) convex polyhedron (cf. [13]), thus the whole information carried by $x$ is determined by a finite amount of data, namely the combinatorics of the faces of $B_{x}$.
Unfortunately, the Thurston norm has been described just in some special families of manifolds (see for example [13, 5, 9, 10, 1, 2]). In [2], we focused on the manifolds obtained as $2$-bridge link complements. There, the Thurston norm is a seminorm on the plane and happens to be always quite simple: the unit polygon has at most $8$ vertices, lying along the axes and the principal bisectors. However, by means of this result, we showed that the Thurston norms obtainable on the plane can have unit balls with an arbitrary number of sides, even if we restrict to complements in $S^{3}$ of links with two unknotted components.
In the present paper, we move our attention to a different family of $3$-manifolds, namely graph manifolds. In many cases, the second homology of a graph manifold has dimension $\geq 3$, a setting where even less is known about the possible shapes of Thurston unit polyhedra. The advantage of graph manifolds is that their surfaces can be easily understood. Indeed, a graph manifold decomposes into Seifert fibered pieces, and an incompressible and boundary-incompressible surface in a Seifert fibered manifold can be isotoped to become either horizontal or vertical. Moreover, a useful result by Neumann [8] helps us to understand when surfaces inside the different pieces can patch together to give a properly embedded surface in the ambient graph manifold.
@@
Theorem 4.1. Let |||||$\cdot$||| | ⋅ | | be a norm on $\operatorname{\mathbb{R}}^{d}$ that can be written as
@@
for some $\beta_{1},...,\beta_{n}\in\operatorname{\mathbb{Q}}^{d}$, and let $g_{1},...,g_{n}$ be nonnegative integers. There exist a closed oriented good graph manifold $M$ with base surfaces of genera $g_{1},...,g_{n}$ and a rational basis on $H_{2}(M;\operatorname{\mathbb{R}})$, such that the nonvanishing Thurston norm on $H_{2}^{nv}(M;\operatorname{\mathbb{R}})$ in the induced basis coincides with |||||$\cdot$||| | ⋅ | |. Moreover, we can ask for either all or none of the top-dimensional Thurston cones of $H_{2}(M;\operatorname{\mathbb{R}})$ to be fibered.
In Theorem 4.1, good graph manifold means that its Seifert fibered JSJ pieces all have orientable base orbifold, whereas the nonvanishing homology $H_{2}^{nv}(M;\operatorname{\mathbb{R}})$ and the nonvanishing Thurston norm refer to the quotient of $H_{2}(M;\operatorname{\mathbb{R}})$ by the null-space of the Thurston norm and the induced norm respectively. Notice that, in the case of good graph manifolds, we can also compute the dimension of the null-space of the Thurston norm.
@@
Observe that, in [13], Thurston has shown a result that is stronger than Corollary 4.4 as it just uses integral bases and gives ways to realise the polygons directly on $H_{2}(M;\operatorname{\mathbb{R}})$, in such a way that any prescribed collection of edges corresponds exactly with fibered cones. At the same time, the nice feature of Corollary 4.4 is that we can keep track of more structure on the manifolds used to realise the given planar norms, namely their being good graph manifolds together with whether they fiber over the circle or not.
@@
Corollary 5.8 Let $P\subset\operatorname{\mathbb{R}}^{n}$ be a polyhedron with rational vertices which is symmetric through the origin. There is a closed oriented good graph manifold $M$, such that there is a rational basis in which $H_{2}^{nv}(M;\operatorname{\mathbb{R}})\cong\operatorname{\mathbb{R}}^{n}$ and the partition of $H_{2}^{nv}(M;\operatorname{\mathbb{R}})$ into cones over the faces of the nonvanishing Thurston unit ball is a refinement of the partition of $\operatorname{\mathbb{R}}^{n}$ into cones over the faces of $\partial P$. Moreover, we can ask for either all or none of the top-dimensional Thurston cones of $H_{2}(M;\operatorname{\mathbb{R}})$ to be fibered.
Structure of the paper. In Section 2 we recall the definitions and basic properties of the main objects on focus: the Thurston norm, Seifert fibered manifolds, and graph manifolds. Along the way, we introduce a result by Neumann [8] that facilitates our understanding of the surfaces sitting inside graph manifolds. In Section 3 we compute the norm of a graph manifold. In Section 4, a characterization of all norms that can be obtained as nonvanishing Thurston norm of graph manifolds is achieved (see Theorem 4.1). Finally, in Section 5 we focus on the polyhedra arising as unit balls of Thurston norms on graph manifolds. This last section aims to understand the geometry of such polyhedra and their properties better. In this way, we will have an idea of how “generic” such polyhedra are, inside the class of all polyhedra or the class of those polyhedra arising as Thurston unit balls of some $3$-manifold.
Aknowledgements. I would like to thank my PhD advisor Mehdi Yazdi for useful discussions on the topic and for his supervision. Many thanks also to the anonymous referee for their thoughtful comments on the initial version of this paper.
@@ 2 Preliminaries
We now recall the basic definitions and properties of the Thurston norm. For further details, we refer to [13]. From now on, $M$ will always be a $3$-dimensional compact orientable smooth manifold.
@@
[13] If $S\subset M$ is the fiber of a fibration of $M$ over $S^{1}$, then $[S]\in H_{2}(M,\partial M)$ lies in the interior of the cone over a top-dimensional boundary face of the Thurston ball $B_{x}$. Viceversa, there is a collection of top-dimensional boundary faces of $B_{x}$ such that every integral class lying in the interior of the cone over one such face is represented by a fiber of a fibration of $M$ over $S^{1}$.
@@
[4, Proposition 11.2] Let $M$ be a compact oriented $3$-manifold with incompressible boundary. Let $T\subset M$ be a disjoint union of incompressible tori and let $M^{\prime}$ be $M$ cut along $T$, seen as a submanifold of $M$. Consider a properly embedded oriented surface $S\subset M$, being transverse to $T$. Indicate with $x_{M}$ and $x_{M^{\prime}}$ the Thurston norms on $H_{2}(M,\partial M)$ and $H_{2}(M^{\prime},\partial M^{\prime})$ respectively. The following equality holds
@@ 2.2 Seifert fibered manifolds and their Thurston norm
We now introduce the building blocks for graph manifolds, which are Seifert fibered manifolds. For further details, see [6] for example.
@@
The manifold $(\Sigma,(p_{1},q_{1}),...,(p_{r},q_{r}))$ can be endowed with a singular fibration by circles (see [6]) over the base orbifold $(\Sigma,p_{1},...,p_{r})$, i.e. the orbifold whose base surface is $\Sigma$ with $r$ cone points, with orders $p_{1},...,p_{r}$. This singular fibration is restricted to the regular circle bundle on $\Sigma\widetilde{\times}S^{1}-\left(\cup_{i}N(c_{i})\right)$ and has singular fibers only corresponding to the cone points of $(\Sigma,p_{1},...,p_{r})$. This special type of singular fibration is called Seifert fibration, hence we call $(\Sigma,(p_{1},q_{1}),...,(p_{r},q_{r}))$ a Seifert fibered manifold (over the base orbifold $(\Sigma,p_{1},...,p_{r})$).
The classification of Seifert fibered manifolds up to homeomorphism is completely understood [6]. A weaker problem concerns the classification of Seifert fibered manifolds up to isomorphism. If we are given two Seifert fibered manifolds $M$ and $M^{\prime}$ with fixed Seifert fibrations on them, we say they are isomorphic if there is a diffeomorphism between $M$ and $M^{\prime}$ respecting the fibrations. The same manifold can admit two different (non-isomorphic) Seifert fibrations.
@@
[6, Theorem 10.4.19] If $\partial\Sigma\neq\emptyset$ and $\Sigma$ is not a disc nor a Moebius band, then $(\Sigma,(p_{1},q_{1}),...,(p_{r},q_{r}))$ admits only one Seifert fibration up to isomorphism.
@@
When dealing with Seifert fibered manifolds, two important invariants are always concerned, namely the base orbifold Euler characteristic and the Euler number. Here, we will use the conventions of [8].
@@
We can think of $\chi^{\text{orb}}(M)$ as a measure of the topological complexity $\chi^{\text{orb}}(\Sigma)$ of the base orbifold. The defining formula is suitably chosen to be multiplicative under finite orbifold coverings, which is also an ingredient used in the proof of Proposition 2.14.
For a closed Seifert fibered manifold $M$, the Euler number is zero exactly when the Seifert fibration is finitely covered by a trivial circle bundle (see [6, Corollary 10.3.29]). We can then consider the Euler number as a measure of the twistedness of the circle fibers in the Seifert fibration.
For the remaining part of this section, let $M=(\Sigma,(p_{1},q_{1}),...,(p_{r},q_{r}))$ with the induced Seifert fibration. The fibration in $M$ helps us to describe the surfaces sitting inside $M$.
@@
Let $S\subset M$ be a properly embedded surface. We say that $S$ is vertical if it is a union of regular circle fibers. We say that $S$ is horizontal if it intersects the regular fibers transversely.
@@
[6, Propositions 10.4.9 and 10.4.10] Let $S\subset M$ be a properly embedded surface and suppose $M$ is not diffeomorphic to $S^{2}\times S^{1}$ nor to $\mathbb{RP}^{2}\widetilde{\times}S^{1}$. If $S$ is incompressible, $\partial$-incompressible and no component of $S$ is $\partial$-parallel, then $S$ is isotopic to either a vertical or a horizontal surface.
@@
Additionally, given any regular fiber $\ell$ of the Seifert fibration, the following formula holds for every $\alpha\in H_{2}(M,\partial M)$:
$x(\alpha)=\chi_{-}^{orb}(M)\cdot|i(\alpha,\ell)|=\chi_{-}^{orb}(M)\cdot|\psi(\alpha)|,$ (1)
where $i(\alpha,\ell)$ denotes the algebraic intersection number, $\psi\in H^{2}(M)\cong\hom(H_{2}(M),\operatorname{\mathbb{R}})$ is the Poincare dual of $\ell$ and $\chi_{-}^{orb}(M)=\max\left(0,-\chi^{orb}(M)\right).$
@@
Let $S\subset M$ be a properly embedded oriented surface. Up to homology, we can assume $S$ is incompressible and $\partial$-incompressible. In particular, after possibly discarding null-homologous components, we can suppose $S$ is either vertical or horizontal.
@@
As proved in [6, Proposition 10.4.8], if $M$ is closed with $e(M)\neq 0$, then $M$ contains no horizontal surface, so $H_{2}^{nv}(M,\partial M)=0$. So, assume that $\partial M\neq\emptyset$ or $e(M)=0$, in which case $M$ admits a horizontal surface.
Suppose that $S$ is horizontal and connected. Then, the projection $M\to\Sigma$ restricts to a degree-$d$ orbifold covering $S\to\Sigma$, where $d$ is the number of times that $S$ transversely intersects each fiber. Thus,
@@
Observe that since $S$ is connected, it must intersect every regular fiber always with the same sign. Indeed, the set of points intersecting the regular fibers positively is open in $S$, as well as the set of points intersecting them negatively, and these two sets together cover $S$. In particular, given any regular fiber $\ell$ of the Seifert fibration,
$d=|S\cap\ell|=|i([S],\ell)|.$ (2)
By cutting $M$ along $S$, we obtain an oriented interval bundle over $S$, hence $S\times[0,1]$. So, the manifold $M$ fibers over the circle with fiber $S$, and this implies that $S$ is norm-minimising in $H_{2}(M,\partial M)$ [13], i.e.
$x([S])=\chi_{-}(S)=d\max(0,-\chi^{orb}(M)).$ (3)
Finally, suppose $S^{\prime}$ is another oriented horizontal surface, transversely intersecting each fiber $d^{\prime}$ times. After possibly reversing the orientation of $S^{\prime}$, we can suppose that the surface given by the oriented cut-and-paste $\overline{S}$ of $d^{\prime}$ parallel copies of $S$ and $d$ parallel copies of $S^{\prime}$ has algebraic intersection 00 with $\ell$. If $\overline{S}$ is homologous to a horizontal surface, such a horizontal surface consists of parallel connected components (since each component is the fiber of a fibration of $M$ over the circle), and thus $i([\overline{S}],\ell)=0$ implies that $\overline{S}$ is null-homologous. If $\overline{S}$ is homologous to a vertical surface, then it annihilates the Thurston norm. This proves that the nonvanishing homology has dimension at most one, the reasoning above shows exactly when it vanishes and formulae (2) and (3) together imply formula (1). ∎
@@ 2.3 Graph manifolds and their reduced plumbing matrices
In this section, we introduce graph manifolds and some results by Neumann [8] that will be essential for our ends. We will be interested only in the closed orientable case, even if the stated facts have broader generality.
@@
When $M$ is a graph manifold, there is a canonical family of decomposing tori in $M$ up to isotopy. This family is characterized as the minimal one (with respect to the inclusion) cutting $M$ into Seifert fibered pieces, and the decomposition is the JSJ decomposition of $M$ [6]. When working with a graph manifold $M$ and unless otherwise mentioned, we will always consider its subdivision into Seifert fibered pieces coming from the JSJ decomposition.
@@
[8, Lemma 4.2] Let $M$ be closed oriented good graph manifold with pieces $M_{1},...,M_{n}$. Assume that $M$ is not a torus-bundle over $S^{1}$. Let $(l_{1},...,l_{n})$ be a tuple of integers. The following properties are equivalent:
  • There are a natural $n_{0}>0$ and a properly embedded oriented surface $S\subset M$ intersecting the regular fibers of $M_{i}$ algebraically $n_{0}l_{i}$ times, for each $i=1,...,n$;
@@
In the original version of [8, Lemma 4.2], in the place of the first property there was
@@
Observe that the corrected version of Lemma 2.17 is enough to prove the results of [8], by just replacing the original lemma.
@@
Sketch of the proof. In [8], Neumann considers $M$ as a plumbing manifold, often relying on the results of [7]. A plumbing manifold is a $3$-manifold $N$ that admits a family of decomposing tori $\mathcal{T}$ such that:
@@
The base surfaces of the plumbing pieces $M_{1}^{\prime},...,M_{k}^{\prime}$ of $M$ give a basis of $H_{2}(X;\operatorname{\mathbb{Z}})\cong\operatorname{\mathbb{Z}}^{k}$. In such basis, $\alpha$ sends the class of a properly embedded oriented surface $S\subset M$ to the tuple $(\ell_{1}^{\prime},...,\ell_{k}^{\prime})$, where $\ell_{i}^{\prime}$ is the algebraic intersection of $S$ with the fiber of $M_{i}^{\prime}$. The map $\beta$ is the intersection form sending a surface $S\subset X$ to the algebraic intersection homomorphism $\langle[S],\cdot\rangle$, represented by a matrix $\mathcal{S}$ in the aforementioned basis. Now, we just need to establish the relation between the tuples $(\ell_{1}^{\prime},...,\ell_{k}^{\prime})$ and $(\ell_{1},...,\ell_{n})$, as well as the relation between $\mathcal{S}$ and the reducing plumbing matrix $A$ of $M$.
Each piece $M_{j}$ of the JSJ decomposition of $M$ corresponds to a subgraph $\Gamma_{j}$ of $\Gamma^{\prime}$, and the subgraphs obtained in this way all together cover $\Gamma^{\prime}$, while pairwise intersecting just in a union of vertices. In every subgraph $\Gamma_{j}$, there is a preferred vertex, called the node of the subgraph. The node is a vertex with positive genus or with degree $\geq 3$. Label the vertices of $\Gamma^{\prime}$ so that the nodes are labelled $1,...,n$. The matrix $\mathcal{S}$ can be modified through a sequence of column and row moves to get the direct sum of $A$ with a diagonal $(k-n)\times(k-n)$ matrix. Each row (resp. column) move consists of adding to a row (resp. column) a rational multiple of one of the last $k-n$ rows (resp. columns). In particular, if $(\ell_{1}^{\prime},...,\ell_{k}^{\prime})\in\operatorname{\mathbb{Z}}^{k}$ annihilates $\mathcal{S}$, then $(\ell_{1}^{\prime},...,\ell_{n}^{\prime})$ annihilates $A$. Viceversa, if $(\ell_{1},...,\ell_{n})\in\operatorname{\mathbb{Z}}^{n}$ annihilates $A$, then there are $\ell_{n+1}^{\prime},...,\ell_{k}^{\prime}\in\operatorname{\mathbb{Q}}$ such that $(\ell_{1},...,\ell_{n},\ell_{n+1}^{\prime},...,\ell_{k}^{\prime})$ annihilates $\mathcal{S}$. Since $\ell_{n+1}^{\prime},...,\ell_{k}^{\prime}$ can be nonintegral, we could need to multiply $(\ell_{1},...,\ell_{n},\ell_{n+1}^{\prime},...,\ell_{k}^{\prime})$ by a nonzero natural number $n_{0}$ to obtain a tuple of integers annihilating $\mathcal{S}$.